simplify each expression for the following
6x +11x =
23ab - (-2ab) =
3x + (-7x) + -10x =
7(3m) - 18 + 2m =
43 - 6(-3e) - 5e =

Found 2 solutions by praseenakos@yahoo.com, malakumar_kos@yahoo.com:
Answer by praseenakos@yahoo.com(507)

Question:

simplify each expression for the following
6x +11x =
23ab - (-2ab) =
3x + (-7x) + -10x =
7(3m) - 18 + 2m =
43 - 6(-3e) - 5e =

Answer


6x +11x = 17x
===================================================
23ab - (-2ab) = 23 ab + 2 ab = 25 ab

====================================================
3x + (-7x) + -10x = 3x - 7x - 10x = -14x
=======================================================

7(3m) - 18 + 2m = 21 m - 18 + 2m = 23 m - 18
[If the above question is 7(3m) - 18m + 2m
[Then the answer is 7(3m) - 18m + 2m = 21 m - 18m + 2m = 5m ]
================================================================
43 - 6(-3e) - 5e = 43 + 18 e - 5e = 43 + 13 e


Hope you understood.

Regards.

Praseena.


Answer by malakumar_kos@yahoo.com(315)


simplify each expression for the following
6x +11x =
23ab - (-2ab) =
3x + (-7x) + -10x =
7(3m) - 18 + 2m =
43 - 6(-3e) - 5e =

1)6x+11x = 17x
2)23ab-(-2ab) = 23ab+2ab = 25ab
3)3x+(-7x)+(-10x) = 3x-7x-10x = x(3-7-10) = -14x
4)7(3m)-18+2m = 21m+2m-18 = 23m - 18
5)43-6(-3e)-5e = 43+18e-5e = 43+13e